For a specific fluorescing compound T, exciting light is absorbed at 280 nm and emitting light is measured at 400 nm.

a) Calculate the light frequency corresponding to each of these wavelengths.

b) Calculate the energy difference between these wavelengths in kJ/mole.
Answer #1

Frequency = velocity of light/wavelength

Energy = Planck's constant*frequency

a) Frequency of 280 nm wave = (3*108)/(280*10-9) = 1.072*1015    hz

b) Frequency of 400 nm wave = (3*108)/(400*10-9) = 7.5*1014    hz

a) Energy of 280 nm wave = (6.63*10-34)*(1.072*1015) = 7.11*10-19 J

energy/mole = energy of 1 wave*Avogadro's number = (7.11*10-19 )*(6.022*1023) = 4.28*105 J/mole = 428 kJ/mole

b) Energy of 400 nm wave = (6.63*10-34)*(7.5*1014) = 4.97*10-19 J

energy/mole = energy of 1 wave*Avogadro's number = (4.97*10-19 )*(6.022*1023) = 2.99*105 J/mole = 299 kJ/mole